a)    Consider an economy characterized by the following relationships:

           Consumption:            C = 20 + 0.5(Y - T)

           Investment:                 I = 25 – 200i

           Money Demand:      Md = Y(0.5 – i)



1
Expert's answer
2021-08-09T06:50:02-0400

"Y=C+I;\nY=20+0.5Y+25-200i;"

"Y=45+0.5Y-200i;"

"0.5Y=45-200i;"

"Y=90-400i;"